# Bell Inequality Violation of Light Quarks in Back-to-Back Dihadron Pair Production at Lepton Colliders

### v1

@@ Abstract
Spin correlations between particles produced at colliders provide valuable insights for quantum information studies. While traditional studies of quantum information at colliders are typically limited to massive particles with perturbative decay, we propose an innovative method to explore the Bell inequality in massless quark pair systems by analyzing the azimuthal correlations in back-to-back $\pi^{+}\pi^{-}$ dihadron pair production at lepton colliders. Revisiting the Belle data, we have shown the potential to detect Bell inequality violation of light quarks by introducing an additional angular cut, achieving a significance of 2.5 $\sigma$ even in the worst-case scenario of 100% correlated systematic uncertainties in each bins. The significance substantially exceeds $5\sigma$ when considering uncorrelated systematic uncertainties. Our approach opens avenues for exploring spin quantum information in the non-perturbative aspect and leverages existing data for quantum information research.
preprint: PITT-PACC-2414
@@
While the decay approach is only valid for massive particle pairs that decay perturbatively, exploring spin entanglement and Bell inequality violations in light quark pairs ($q\bar{q}$) presents a compelling research avenue that has been overlooked in the literature. In contrast to the top quark, the light quark does not decay but instead fragments into a jet of hadrons after produced from hard scattering. Nevertheless, the spin of a light quark can still be transferred to its hadron final states and can be systematically characterized by the non-perturbative fragmentation functions. Within the framework of collinear factorization, it has been demonstrated that only the transverse spin of a light quark can be transferred to the unpolarized hadron final states, and reconstructing it requires the presence of at least two hadrons $h_{a}(\vec{p}_{a})$ and $h_{b}(\vec{p}_{b})$ within the jet [27]. The transverse spin $\vec{s}_{T}$ of a quark disrupts the rotational symmetry around its momentum direction, leading to an azimuthal asymmetry in the jet constituents. This asymmetry is not accounted for in perturbative calculations but captured by interference dihadron fragmentation functions (diFFs) [27, 28, 29, 30, 31, 32, 33]  [^1] The linear polarization of gluons from initial states could result in a similar azimuthal asymmetry [34].. Thus, the back-to-back dihadron pair production $e^{+}e^{-}\to q\bar{q}\to(h_{a1},h_{b1})+(h_{a2},h_{b2})+X$ offers a promising avenue for investigating the transverse spin correlation of $q\bar{q}$ produced at lepton colliders [35, 36].
@@
The massless quark pair system also manifests a significant advantage for Bell inequality studies, as the transverse spins of $q$ and $\bar{q}$ generated from the Drell-Yan processes are 100% correlated in the central scattering region. The corresponding spin state forms a maximally entangled Bell state $\frac{1}{\sqrt{2}}(\ket{\uparrow_{y}\downarrow_{y}}+\ket{\downarrow_{y}\uparrow_{y}})$, with |$\ket{\uparrow_{y}}$ and |$\ket{\downarrow_{y}}$ denoting the transverse polarized eigenstates along $\hat{y}$, the normal direction of the hard scattering plane. Therefore, although the complete $q\bar{q}$ spin density matrix, encompassing longitudinal spin, cannot be reconstructed from back-to-back dihadron pairs in the collinear limit, the transverse spin already presents the most relevant information about the $q\bar{q}$ spin Bell state. By applying a simple selection cut on the hard scattering angle, an ideal ensemble of the spin Bell state can be prepared from the Drell-Yan process. The maximal entangled behavior makes the $q\bar{q}$ system attractive for investigating quantum information, especially, offering the potential for a distinct signal of Bell inequality violation that has not yet been observed at colliders using the decay approach.
@@
In this Letter, we propose a novel approach to perform a partial quantum tomography of $q\bar{q}$ spin state by utilizing dihadron interference fragmentation functions as spin analyzing powers to reconstruct the transverse parts in the $q\bar{q}$ spin density matrix, and present a sufficient condition for detecting entanglement and Bell inequality violation in $q\bar{q}$ production at lepton colliders. Our approach represents the first exploration of Bell inequality violation within the massless quark pair system, opening up avenues for spin quantum information research in a non-perturbative perspective.
@@
With the method we have developed, numerous existing data becomes readily available for quantum information studies. In this Letter we revisit the back-to-back $\pi^{+}\pi^{-}$ dihadron pair production at the $10.58$ GeV Belle collider [37]. Our analysis indicates that the current uncertainty implies a potential observation of the Bell inequality violation in the $q\bar{q}$ system from existing data, with a significance of at least $2.5\,\sigma$. A more detailed examination of systematic uncertainties is expected to yield an even much larger significance, possibly resulting in the first measurement of Bell inequality violation in a collider environment.
@@
where the Bell variable $\mathcal{B}$ can be either $\mathcal{B}_{+}$ or $\mathcal{B}_{-}$. The maximal value of the Bell variable in quantum theory is $\mathcal{B}=2$, which is achieved when the state is a maximally entangled Bell state.
In the Standard Model (SM), the spin correlation matrix of $q\bar{q}$ in the Drell-Yan process $e^{-}e^{+}\to q\bar{q}$ at Belle is
@@
where $\Theta$ represents the hard scattering angle, and the net polarization vectors of $q$ and $\bar{q}$ are zero, i.e., $B_{i}=\bar{B}_{i}=0$. Then the Bell variables are
@@
where $\rho(\Theta)$ is the density matrix of the $q\bar{q}$ produced at a fixed scattering angle $\Theta$, and $\sigma_{\rm hard}$ is the hard scattering cross section within the angular cut. The angular-averaged state is also known as “fictitious state” [5] due to introducing the subtleties of basis dependence. However, for massless quarks, the conventional choice of helicity basis is already optimal for testing Bell inequality violation [40, 41].
@@
Fragmentation of quark pair spin state.—We consider the Drell-Yan process $e^{+}e^{-}\to q(k_{1})\bar{q}(k_{2})\to(\pi^{+}\pi^{-})+(\pi^{+}\pi^{-})+X$, with the leading-order kinematics depicted in Fig. 1. The kinematics of a $\pi^{+}\pi^{-}$ pair are characterized by their total momentum $P_{i}=p_{i}^{\pi^{+}}+p_{i}^{\pi^{-}}$ (for $i=1,2$) and their momentum difference $R_{i}\equiv(p_{i}^{\pi^{+}}-p_{i}^{\pi^{-}})/2$. We consider the collinear limit where the total momentum $\vec{P}_{i}$ is aligned with the quark momentum $\vec{k}_{i}$ and approximately the jet directions, and the dihadron fragmentation phase space in this limit is conventionally described by the dihadron invariant mass $M_{i}\equiv P_{i}^{2}$, the azimuthal angle $\phi_{i}$ of the relative transverse momentum $\vec{R}_{i,T}$, and the total momentum fractions $z_{1}=P_{1}^{-}/k_{1}^{-}$ and $z_{2}=P_{2}^{+}/k_{2}^{+}$ where $P_{i}^{\pm}(k_{i}^{\pm})$ represent the light-cone components of the momentum of $P_{i}(k_{i})$.
Under the collinear factorization, the differential distribution of back-to-back $\pi^{+}\pi^{-}$ pairs fragmented from a quark pair with spin density matrix $\rho_{ss^{\prime},\bar{s}\bar{s}^{\prime}}$ can be written as [28]:
$\frac{\mathrm{d}\sigma}{\mathrm{d}\Omega_{1}\mathrm{d}\Omega_{2}}=\sigma_{\rm hard}\sum_{ss^{\prime}\bar{s}\bar{s}^{\prime}}\rho_{ss^{\prime},\bar{s}\bar{s}^{\prime}}\mathcal{D}^{ss^{\prime}}_{\pi^{+}\pi^{-}/q}\mathcal{D}^{\bar{s}\bar{s}^{\prime}}_{\pi^{+}\pi^{-}/\bar{q}},$ (9)
where $\mathrm{d}\Omega_{i}=\mathrm{d}z_{i}\mathrm{d}M_{i}\mathrm{d}\phi_{i}$ represents the phase space of the $\pi^{+}\pi^{-}$ pair fragmented from $q$ or $\bar{q}$, $(s,s^{\prime},\bar{s},\bar{s}^{\prime})$ are spin indices of $q$ and $\bar{q}$, $\mathcal{D}_{\pi^{+}\pi^{-}/q}^{ss^{\prime}}$ represents the dihadron fragmentation matrix of the parton $q$ into a $\pi^{+}\pi^{-}$ pair, and $\sigma_{\rm hard}$ is the hard scattering cross section. The scattering process is therefore factorized into two components: the hard parts $\sigma_{\rm hard}$ and $\rho_{ss^{\prime},\bar{s}\bar{s}^{\prime}}$ that can be calculated perturbatively, e.g., the SM predictions in Eq. (5), and the collinear fragmentation part $\mathcal{D}_{\pi^{+}\pi^{-}/q}$ that can only be determined through global fitting with experimental data. For unpolarized final-state hadrons, the parity reflection symmetry results in only two independent collinear factors for each flavor of quarks:
@@
where ±$\mathcal{B}_{\pm}=C_{xx}\pm C_{yy}$ are the Bell variables defined in Eq.(3), and $e_{q}$ denotes the electric charge of the quark $q$. Note that $CP$ symmetry is assumed so that there is no sine terms in the distribution. With the SM predictions of $\mathcal{B}_{\pm}$ in Eq. (6), Eq. (Bell Inequality Violation of Light Quarks in Back-to-Back Dihadron Pair Production at Lepton Colliders) reproduces the well-known Artru-Collins asymmetry [35, 36].
@@
Thus, the Bell variable is reconstructed from two parts, the experimental asymmetry observable $A_{12}$ that depends on the strength of transverse spin correlation of $q\bar{q}$, and the theory prediction $\alpha^{z_{1},z_{2}}_{M_{1},M_{2}}$ that is a universal spin analyzing power independent of the $q\bar{q}$ spin state.
@@
Since the covariance matrix is not explicitly provided in Ref. [37], we assume the systematic uncertainties are either 0% or 100% correlated, while both the statistical uncertainties and the theory uncertainties are considered uncorrelated. The expected results of $\mathcal{B}_{-}$ under both assumptions are depicted in Fig. 3. Given the significant number of events collected in Belle ($\sim 10^{8}$), the statistical uncertainty is much smaller than the systematic uncertainty. Consequently, imposing a more stringent cut does not have much effect on the uncertainty, but yields a purer Bell state with a larger signal of Bell inequality violation. By selecting an angular cut of $|\cos\Theta|<0.1$ above the jet resolution [44], one can analyze a highly pure Bell state with $\text{Tr}(\rho^{2})>0.99$ and anticipate detecting Bell inequality violation with a sensitivity of 2.5 $\sigma$, even in the worst-case scenario where the systematic uncertainties in each bin are 100% correlated. In the case of assuming uncorrelated systematic uncertainties, we expect to achieve a sensitivity of 6.7 $\sigma$.
@@
Conclusions and Discussions.—We propose a novel approach to investigate the Bell inequality violation of the massless quark pairs by utilizing the interference diFF in back-to-back $\pi^{+}\pi^{-}$ dihadron production at lepton colliders. We demonstrated that the Bell inequality of quark pair system can be measured with a single observable, the azimuthal angle correlation of the back-to-back dihadron pairs. Our method represents the first examination of Bell inequality violation in massless quark pairs and the first exploration of a highly pure spin Bell state produced at colliders. By implementing a straightforward additional selection cut on the hard scattering angle, we demonstrated that the measurement of Bell inequality violation can achieve a significance of 2.5 $\sigma$ in the worst-case scenario of 100% correlated systematic uncertainties for the current Belle experiment, while the significance substantially exceeds $5\sigma$ when considering uncorrelated systematic uncertainties. Therefore, by comprehensively addressing systematic uncertainties, it becomes feasible to confirm the violation of the Bell inequality using available data. The detection of Bell inequality violation would also imply the presence of entanglement, even though direct measurement of quantum concurrence is not achievable in this process.
@@
While we examine dihadron pair production within the collinear factorization framework as a first example to investigate the Bell inequality violation of the massless quark pairs, our method is equally applicable to other fragmentation channels within the TMD framework. Combining various channels in the future would help reduce both theoretical and experimental uncertainties associated with studying these quantum phenomena.
@@ Acknowledgements.
We thank Chris Cocuzza for the help on the usage of JAMDiFF, and thank Matthew Low and Zhite Yu for the helpful discussion. Kun Cheng is supported in part by U.S. Department of Energy under grant No. DE-SC0007914. Bin Yan is supported in part by the National Science Foundation of China under Grant No. 12422506, the IHEP under Grant No. E25153U1 and CAS under Grant No. E429A6M1.

### v2

# Bell Inequality Violation of Light Quarks in Dihadron Pair Production at Lepton Colliders

@@ Abstract
Spin correlations between particles produced at colliders provide valuable insights for quantum information studies. While traditional studies of quantum information at colliders are typically limited to massive particles with perturbative decay, we propose an innovative method to explore the Bell inequality in massless quark pair systems by analyzing the azimuthal correlations in $\pi^{+}\pi^{-}$ dihadron pair production at lepton colliders. Revisiting the Belle data, we have shown the potential to detect Bell inequality violation of light quarks by introducing an additional angular cut, achieving a significance of 2.5 $\sigma$ even in the worst-case scenario of 100% correlated systematic uncertainties in each bins. The significance substantially exceeds $5\sigma$ when considering uncorrelated systematic uncertainties. Our approach opens avenues for exploring spin quantum information with non-perturbative processes as spin analyzer and leverages existing data for quantum information research.
preprint: PITT-PACC-2414, CPTNP-2025-020
@@
While the decay approach is only valid for massive particle pairs that decay perturbatively, exploring spin entanglement and Bell inequality violations in light quark pairs ($q\bar{q}$) presents a compelling research avenue that has been overlooked in the literature. In contrast to the top quark, the light quark does not decay but instead fragments into a jet of hadrons after produced from hard scattering. Nevertheless, the spin of a light quark can still be transferred to its hadron final states and can be systematically characterized by the non-perturbative fragmentation functions. Within the framework of collinear factorization, it has been demonstrated that only the transverse spin of a light quark can be transferred to the unpolarized hadron final states, and reconstructing it requires the presence of at least two hadrons $h_{a}(\vec{p}_{a})$ and $h_{b}(\vec{p}_{b})$ within the jet [27]. The transverse spin $\vec{s}_{T}$ of a quark disrupts the rotational symmetry around its momentum direction, leading to an azimuthal asymmetry in the jet constituents. This asymmetry is not accounted for in perturbative calculations but captured by interference dihadron fragmentation functions (diFFs) [27, 28, 29, 30, 31, 32, 33, 34]  [^1] The linear polarization of gluons could result in a similar azimuthal asymmetry [35, 36].. Thus, the dihadron pair production $e^{+}e^{-}\to q\bar{q}\to(h_{a1},h_{b1})+(h_{a2},h_{b2})+X$ offers a promising avenue for investigating the transverse spin correlation of $q\bar{q}$ produced at lepton colliders [37, 38].
@@
The massless quark pair system also manifests a significant advantage for Bell inequality studies, as the transverse spins of $q$ and $\bar{q}$ generated from the $e^{+}e^{-}\to\gamma^{*}\to q\bar{q}$ processes are 100% correlated in the central scattering region. The corresponding spin state forms a maximally entangled Bell state $\frac{1}{\sqrt{2}}(\ket{\uparrow_{y}\downarrow_{y}}+\ket{\downarrow_{y}\uparrow_{y}})$, with |$\ket{\uparrow_{y}}$ and |$\ket{\downarrow_{y}}$ denoting the transverse polarized eigenstates along $\hat{y}$, the normal direction of the hard scattering plane. Therefore, although the complete $q\bar{q}$ spin density matrix, encompassing longitudinal spin, cannot be reconstructed from dihadron pairs in the collinear limit, the transverse spin already presents the most relevant information about the $q\bar{q}$ spin Bell state. By applying a simple selection cut on the hard scattering angle, an ideal ensemble of the spin Bell state can be prepared from the $e^{+}e^{-}$ annihilation. The maximal entangled behavior makes the $q\bar{q}$ system attractive for investigating quantum information, especially, offering the potential for a distinct signal of Bell inequality violation that has not yet been observed at colliders using the decay approach.
@@
In this Letter, we propose a novel approach to perform a partial quantum tomography of $q\bar{q}$ spin state by utilizing dihadron interference fragmentation functions as spin analyzing powers to reconstruct the transverse parts in the $q\bar{q}$ spin density matrix, and present a sufficient condition for detecting entanglement and Bell inequality violation in $q\bar{q}$ production at lepton colliders. Our approach represents the first exploration of Bell inequality violation within the massless quark pair system, opening up avenues for spin quantum information research with non-perturbative processes as spin analyzer.
@@
With the method we have developed, numerous existing data becomes readily available for quantum information studies. As a first example, we revisit the $\pi^{+}\pi^{-}$ dihadron pair production at the $10.58$ GeV Belle collider [39]. Our analysis indicates that the current uncertainty implies a potential observation of the Bell inequality violation in the $q\bar{q}$ system from existing data, with a significance of at least $2.5\,\sigma$. A more detailed examination of systematic uncertainties is expected to yield an even much larger significance, possibly resulting in the first measurement of Bell inequality violation in a collider scattering process.
@@
where the Bell variable $\mathcal{B}$ can be either $\mathcal{B}_{+}$ or $\mathcal{B}_{-}$. The maximal value of the Bell variable in quantum theory is $\mathcal{B}=2$, which is achieved when the state is a maximally entangled Bell state. In the following we only focus on the transverse parts $C_{xx}$ and $C_{yy}$ of the spin correlation of $q\bar{q}$ as they are sufficient to measure Bell inequality violation.
In the Standard Model (SM), the spin correlation matrix of $q\bar{q}$ in the $e^{-}e^{+}\to q\bar{q}$ process at Belle is
@@
where $\Theta$ represents the hard scattering angle, and the net polarization vectors of $q$ and $\bar{q}$ are zero, i.e., $B_{i}=\bar{B}_{i}=0$. We see that the transverse spin of $q$ and $\bar{q}$ along $x$($y$) direction are positively(negatively) correlated [42], both achieving 100% correlation when $\Theta=\pi/2$. Then the Bell variables are
@@
where $\rho(\Theta)$ is the density matrix of the $q\bar{q}$ produced at a fixed scattering angle $\Theta$ with azimuthal angle of $q\bar{q}$ averaged over, and $\sigma_{\rm hard}$ is the hard scattering cross section within the angular cut. The angular-averaged state is also known as “fictitious state” [4] due to introducing the subtleties of basis dependence. However, for massless quarks, the conventional choice of helicity basis is already optimal for testing Bell inequality violation [43, 44].
@@
Fragmentation of quark pair spin state.—We consider the semi-inclusive process $e^{+}e^{-}\to q(k_{1})\bar{q}(k_{2})\to(\pi^{+}\pi^{-})+(\pi^{+}\pi^{-})+X$, with the leading-order kinematics depicted in Fig. 1. The kinematics of a $\pi^{+}\pi^{-}$ pair are characterized by their total momentum $P_{i}=p_{i}^{\pi^{+}}+p_{i}^{\pi^{-}}$ (for $i=1,2$) and their momentum difference $R_{i}\equiv(p_{i}^{\pi^{+}}-p_{i}^{\pi^{-}})/2$. We consider the collinear limit where the total momentum $\vec{P}_{i}$ is aligned with the quark momentum $\vec{k}_{i}$ and approximately the jet directions, and the dihadron fragmentation phase space in this limit is conventionally described by the dihadron invariant mass $M_{i}\equiv P_{i}^{2}$, the azimuthal angle $\phi_{i}$ of the relative transverse momentum $\vec{R}_{i,T}$, and the total momentum fractions $z_{1}=P_{1}^{+}/k_{1}^{+}$ and $z_{2}=P_{2}^{-}/k_{2}^{-}$ where $P_{i}^{\pm}(k_{i}^{\pm})$ represent the light-cone components of the momentum of $P_{i}(k_{i})$.
Under the collinear factorization ($|\mathbf{P}_{i}|\gg M_{i}$), the differential distribution of $\pi^{+}\pi^{-}$ pairs fragmented from a quark pair with spin density matrix $\bar{\rho}_{ss^{\prime},\bar{s}\bar{s}^{\prime}}$ can be written as [45, 28, 46]:
$\frac{\mathrm{d}\sigma}{\mathrm{d}\Omega_{1}\mathrm{d}\Omega_{2}}=\sigma_{\rm hard}\sum_{ss^{\prime}\bar{s}\bar{s}^{\prime}}\bar{\rho}_{ss^{\prime},\bar{s}\bar{s}^{\prime}}\mathcal{D}^{ss^{\prime}}_{\pi^{+}\pi^{-}/q}\mathcal{D}^{\bar{s}\bar{s}^{\prime}}_{\pi^{+}\pi^{-}/\bar{q}},$ (9)
where $\mathrm{d}\Omega_{i}=\mathrm{d}z_{i}\mathrm{d}M_{i}\mathrm{d}\phi_{i}$ represents the phase space of the $\pi^{+}\pi^{-}$ pair fragmented from $q$ or $\bar{q}$, $(s,s^{\prime},\bar{s},\bar{s}^{\prime})$ are spin indices of $q$ and $\bar{q}$, $\mathcal{D}_{\pi^{+}\pi^{-}/q}^{ss^{\prime}}$ represents the dihadron fragmentation matrix of the parton $q$ into a $\pi^{+}\pi^{-}$ pair, and $\sigma_{\rm hard}$ is the hard scattering cross section within the angular selection. The scattering process is therefore factorized into two components: the hard parts $\sigma_{\rm hard}$ and $\bar{\rho}_{ss^{\prime},\bar{s}\bar{s}^{\prime}}$ that can be calculated perturbatively, e.g., the SM predictions in Eq. (5), and the collinear fragmentation part $\mathcal{D}_{\pi^{+}\pi^{-}/q}$ that can only be determined from global fittings with experimental data. For unpolarized final-state hadrons, the parity reflection symmetry results in only two independent collinear factors for each flavor of quarks:
@@
where ±$\mathcal{B}_{\pm}=C_{xx}\pm C_{yy}$ are the Bell variables defined in Eq.(3), and $e_{q}$ denotes the electric charge of the quark $q$. Note that $CP$ symmetry is assumed so that there is no sine terms in the distribution. With the SM predictions of $\mathcal{B}_{\pm}$ in Eq. (6), Eq. (Bell Inequality Violation of Light Quarks in Dihadron Pair Production at Lepton Colliders) reproduces the well-known Artru-Collins asymmetry [37, 38].
@@
Thus, the Bell variable is reconstructed from two parts, the experimental asymmetry observable $A_{12}$ that depends on the strength of transverse spin correlation of $q\bar{q}$, and the theory prediction of the spin analyzing power $\alpha^{z_{1},z_{2}}_{M_{1},M_{2}}$ that is independent of the $q\bar{q}$ spin state.
@@
Since the covariance matrix is not explicitly provided in Ref. [39], we assume the systematic uncertainties are either 0% or 100% correlated, while both the statistical uncertainties and the theory uncertainties are considered uncorrelated. The expected results of $\mathcal{B}_{-}$ under both assumptions are depicted in Fig. 3. Given the significant number of events collected in Belle ($\sim 10^{8}$), the statistical uncertainty is much smaller than the systematic uncertainty. Consequently, imposing a more stringent cut does not have much effect on the uncertainty, but yields a purer Bell state with a larger signal of Bell inequality violation. By selecting an angular cut of $|\cos\Theta|<0.1$ above the jet resolution [49], one can analyze a highly pure Bell state with $\text{Tr}(\bar{\rho}^{2})>0.99$ and anticipate detecting Bell inequality violation with a sensitivity of 2.5 $\sigma$, even in the worst-case scenario where the systematic uncertainties in each bin are 100% correlated. In the case of assuming uncorrelated systematic uncertainties, we expect to achieve a sensitivity of 6.2 $\sigma$.
Conclusions and Discussions.—The collider environment is appealing for quantum information studies with many highly demanding quantum measurements naturally implemented, inspiring both fields of high-energy physics and quantum information [50, 51, 15, 52, 53, 54, 55, 18, 56]. Of the rich quantum systems produced at colliders, previous studies are mostly limited to massive particles with perturbative decay. Establishing the framework of studying quantum information for light quarks would make vast existing data available and allow us to make better use of the potential of colliders for quantum information study.
@@
We propose a novel approach to investigate the Bell inequality violation of the massless quark pairs by utilizing the interference diFF in $\pi^{+}\pi^{-}$ dihadron pair production at lepton colliders. We demonstrated that the Bell inequality of quark pair system can be measured with a single observable, the azimuthal angle correlation of the dihadron pairs. Our method represents the first examination of Bell inequality violation in massless quark pairs and the first exploration of a highly pure spin Bell state produced at colliders. By implementing a straightforward additional selection cut on the hard scattering angle, we demonstrated that the measurement of Bell inequality violation can achieve a significance of 2.5 $\sigma$ in the worst-case scenario of 100% correlated systematic uncertainties for the current Belle experiment, while the significance substantially exceeds $5\sigma$ when considering uncorrelated systematic uncertainties. Therefore, by comprehensively addressing systematic uncertainties, it becomes feasible to confirm the violation of the Bell inequality using available data, possibly resulting in the first measurement of Bell inequality violation in a collider scattering process. The detection of Bell inequality violation would also imply the presence of entanglement.
@@
We examine dihadron pair production within the collinear factorization framework as a first example to investigate the Bell inequality violation of the massless quark pairs. With a more careful treatment, it is also possible to generalize our method within the TMD framework [57, 30, 58], which we leave as future works
@@ Acknowledgements.
We thank Chris Cocuzza for the help on the usage of JAMDiFF, and thank Matthew Low, Xin-Kai Wen and Zhite Yu for the helpful discussion. Kun Cheng is supported in part by U.S. Department of Energy under grant No. DE-SC0007914. Bin Yan is supported in part by the National Science Foundation of China under Grant No. 12422506, the IHEP under Grant No. E25153U1 and CAS under Grant No. E429A6M1. The authors gratefully acknowledge the valuable discussions and insights provided by the members of the Collaboration on Precision Tests and New Physics (CPTNP).